OpenStudy (gdvaa67):

Archaeologists find an ancient plaza in Peru. The plaza is shaped like a circle with an area of 1,662 square feet. What is the circumference of the plaza to the nearest foot? 145 feet 135 feet 125 feet 115 feet

OpenStudy (mikey):

do they give you any type of equations

OpenStudy (gdvaa67):

nope

OpenStudy (3mar):

first, we find the radius of the circle; \[area of circle=\pi(r)^2\]

OpenStudy (gdvaa67):

how do we do that exactly?

OpenStudy (3mar):

\[1662=\frac{ 22 }{ 7 }*r^2\] \[r^2=529.031\] \[r=23\] ft

OpenStudy (3mar):

next step is to find the perimeter (cirfcumference) of that plaxza: \[2\pi r=2*\frac{ 22 }{ 7 }*23=144.5175\] ft

OpenStudy (gdvaa67):

so its 145 feet?

OpenStudy (3mar):

plus 0.5175 :) haha

OpenStudy (gdvaa67):

ohh haaha lol

OpenStudy (3mar):

Thanks for the medal! I really appreciate that!

OpenStudy (gdvaa67):

No problem :) you did great thank you for helping me

OpenStudy (3mar):

You are welcome! Any time.

OpenStudy (gdvaa67):

:)

OpenStudy (gdvaa67):

so 145.035

OpenStudy (3mar):

Yes

OpenStudy (3mar):

If you have any problem else, I am ready, Won't be late for you,